OpenStudy (anonymous):

Janet drove 20 miles south and then 30 miles west. How far was she from her original starting point?

OpenStudy (kinggeorge):

This will require use of the pythagorean theorem. Can you tell me what the pythagorean theorem says?

OpenStudy (anonymous):

x^2+x^2=x

OpenStudy (kinggeorge):

Kind of. You're on the right track. It says that if you have a right triangle, and your two legs have length \(a\) and \(b\), then the hypotenuse will have length \[\sqrt{a^2+b^2}.\]Now we just need to convert your question into something where we can use this fact.

OpenStudy (kinggeorge):

This diagram roughly shows how Janet drove. Can you see how we can use the pythagorean theorem?

OpenStudy (anonymous):

Yea, if it was a triangle then that missing part would be an x. So it would be 20^2+30^2=x?

OpenStudy (kinggeorge):

Right, you just need to put it in a square root. So \[x=\sqrt{20^2+30^2}\]

OpenStudy (anonymous):

It'll be 1300, but that won't equal a perfect root if i root it. So should I use the estimate of 36.056 or